What is the volume of 3.7 mol of an ideal gas
at a pressure of 3 atm and a temperature of
0 ◦ C? 1 liter = 0.001 m3 and 1 atm = 101300 Pascals.
Answer in units of L.
1
Expert's answer
2020-03-11T14:18:30-0400

The equation of state for an ideal gas says


"PV=nRT."

Hence, the volume of a gas


"V=\frac{nRT}{P}=\frac{3.7\times 8.3\times 273}{3\times 101300}=0.0276\:\rm m^3=27.6\:\rm L."